Solve the following equation. 6 12 X-1 x2 - 1
Answer #1


Solutions viven, 6 12 2c²f = x- Using cross-multiplication, 667-1) = 12 (x-1) [a²- 6² = 6th lab G{(x-1)(x +1) } = 12 (x-1) &